# LLM-based Property-based Test Generation for Guardrailing Cyber-Physical Systems

### v1

@@ Abstract
Index Terms:
Property-based Testing; LLM4SE; Cyber-Physical System; Safety
I Introduction
In this paper, we propose a novel automated and scalable approach for guardrailing cyber-physical systems (CPS) using property-based test (PBT) generated by Large Language Models (LLMs). Our approach benefits from two major established facts in software engineering: 1) the cyber side of CPSs is essentially a software program amenable to existing automated program analysis tools; and 2) advanced LLMs are strong in analyzing programs and extracting their expected properties [18]. Based on these two observation, our proposed approach, called ChekProp, uses LLMs to generate property-based tests for CPSs before their deployment, i.e., at design time. These PBTs can then be utilized after deployment, i.e., at run time, to detect unsafe behavior of the CPS. ChekProp uses the source code, documentation, and unit tests of the target CPS to extract properties regarding its expected behavior. ChekProp also generates PBTs that verify that the extracted properties hold for the CPS. We implement a prototype of ChekProp and make it publicly available to the community [4].
@@ I Introduction
II Background on Property-based Testing
@@ III-B Initial PBT Generation
Note that, as explained in Section II, there is significant difference between unit tests and property based tests. The unit test only checks the behavior of the program for a specific input. For example, in the unit test (Section 3) of Figure 3, specific total_time, cylinder_interval, etc. are used. Also, in unit testing we usually check that the output is exactly what is expected according to an oracle [17]. In contrast, property based tests check that a more general condition is meet by the program behavior over a wide range of inputs.
The fourth and final part of the Figure 3 instructs the LLM to generate the desired PBTs.
@@
At the end of the initial PBT generation step, ChekProp obtains a set of initial PBTs. ChekProp runs these PBTs and collects their results using an analyzer unit. If a group of generated PBTs fails, ChekProp collects their failure message and enters its PBT improvement loop step as described in Section III-C.
III-C PBT Improvement Loop
@@ III-E Implementation
III-F Property-based Monitoring
@@
More generally, as explained in Section II and as seen in 2, the PBTs generated in the first phase consist of three components: an input generator, a test body, and a test assertion. In the property-based monitoring phase, the PBT input generator is no longer needed, as the inputs are generated by the cyber part (in the form of control commands) and the physical system (via sensor values). The test body is replaced by the monitor that sits between the controller and the physical system collecting relevant data. The properties derived at design time are transformed into guards that are checked at runtime. The monitor verifies that these guards hold in the current state of the system. Based on the situation, if a violation is detected, the monitor can intercept and block the command being sent to the physical system. In this way, the PBTs generated by ChekProp serve as runtime guardrails for CPSs.
IV Experiments
IV-A Research Questions
@@
IV-B Dataset
Table I: The cyber-physical system Python programs considered in our dataset.
@@ IV-B Dataset
P1 Temperature Control System (TCS) Model-based Example [10] The TCS system controls the temperature of a room. The room has a window that blows cold or hot wind into the room. A sensor collects the current temperature and a HC unit is used to heat up or cool down the room environment. Finally, a controller collects the temperature data from the sensor and decides if the HC unit should be used.
@@
P2 Pneumatic Control System (PCS) Model-based Example [10] The PCS regulates the movement of components of a mechanical system that consist of a horizontal and a vertical cylinder, their corresponding controllers and sensors. This mechanical system is designed to pick an object from the ground and move it to a new place.
@@
P3 Laser Tripwire Raspberry Pi Projects [13] A prototype on a breadboard to detect whether a beam of light is hitting the light-dependent resistor (LDR).
@@
P4 Line Following Robot Raspberry Pi Projects [13] A robot with two motors and two sensors. The robot uses its right and left sensors to follow a line and decide if should go turn right, left, or go forward.
@@
P5 Ultrasonic Theremin Raspberry Pi Projects [13] A system that produces a sound with a volume level corresponding to the user distance from the device.
@@
P6 Remote Buggy Raspberry Pi Projects [13] A robot that moves in four directions and a user remotely controls it.
@@
P7 Quick Reaction Game Raspberry Pi Projects [13] A game in which a light turns on and off and two players compete to hit their button faster after the light is turned off.
@@
P8 Presence Indicator Raspberry Pi Projects [13] A monitor made of a list of LEDs that shows the number of people present at a place.
@@
P9 gpiozero InputDevice gpiozero Library Class [15] A core class in gpiozero library that handles connection to input devices on a Raspberry Pi board.
@@
As the current version of ChekProp supports PBT generation for Python programs, we curate a dataset of Python CPS programs for our experiments. This dataset consists of nine programs presented in Table I. These programs are taken from three main sources as follows.
@@
First, we include the Python version of two CPSs that are widely studied in the model checking literature [10]: a temperature control system (TCS) and a pneumatic control system (PCS) (P1 and P2 in Table I). TCS and PCS are presented with the IDs P1 and P2 in Table I. Moradi et al. [10] use the model checking tool of Rebeca  [^1] Rebeca is an actor-based modeling language, in which actors are the units of concurrency and communicate via asynchronous message passing [14]., Afra [14], to detect potential attacks against these systems. For this, they define the correctness properties for each system and assess if Afra can find counterexamples for these properties on models augmented with malicious behavior. The manually defined properties in [10] set a ground-truth with which we can compare the properties automatically extracted by ChekProp.
@@
The second set of programs in our dataset are the six programs (P3-P8 in Table I) taken from open-source Raspberry Pi projects [13] that use the gpiozero library [15]. gpiozero is a Python library that real-world cyber-physical systems employ to connect to Raspberry Pi boards. We take the source of these programs from the official Raspberry Pi website [13] and manually add unit tests for them. These unit tests can be used in ChekProp prompts (see Section III-A). The six programs in our dataset that use gpiozero enable us to evaluate ChekProp applicability on projects that adopt the widely used Raspberry Pi boards.
@@ 69 If :data:‘True‘, the device uses a pull-up resistor to set the GPIO pin
The last program considered in our dataset (P9 in Table I) is InputDevice, a core class from the gpiozero library. This class provides an interface for Python programs to interact with Raspberry Pi input devices, such as barometers, temperature sensors, etc. The InputDevice class is a complex class from gpiozero that connects to various components in this library. Testing this class requires a detailed understanding of the inner workings of the library and how it represents and handles the physical environment. For example, to instantiate an object of the InputDevice class, we have to pass the number of a pin to its constructor method. The type of this pin should be “input”, while some of the pins on a Raspberry Pi board are reserved only for “output”. A correct test should use a pin of the correct type to instantiate an InputDevice. Another example of such details in gpiozero is how pins are activated. To activate a pin on a Raspberry Pi board, its voltage should go high, which can happen by calling the pin.drive_high() method. Writing a test for a CPS program requires an accurate understanding of how interacting with the program, such as calling pin.drive_high(), impacts the physical status of the system, such as increasing the voltage on the Raspberry Pi board. To evaluate whether ChekProp generated PBTs capture such details about CPS programs, we include the InputDevice class in our dataset.
4 presents the InputDevice class. This class, similar to other gpiozero classes, has well-written documentation (lines #2-39). We consider this documentation as the natural language description in ChekProp prompts (see Section III-A). Moreover, gpiozero has extensive unit tests for its classes, which we use them to produce our initial prompt . This confirms that gpiozero classes have the essential components for applying ChekProp: the natural language description, the source code, and the unit test (see Section III-A).
@@
IV-C RQ1: Property Relevance
IV-C1 Methodology
To assess the quality of extracted properties, we run our approach on the nine programs in our dataset and analyze the relevance of the extracted properties. In this experiment, we abstract away the implementation details of the generated PBTs; instead, we focus on how well the properties considered in the PBTs validate the logic of the program under test. For this purpose, we compare properties extracted by our approach with manually crafted properties that we consider as ground-truth. In particular, we evaluate whether the logic checked by ground-truth properties is also validated by ChekProp extracted properties and vice versa. As explained in Section IV-B, for programs P1 and P2, the ground-truth properties are already stated by Moradi et al. [10]. For the remaining programs (P3-P9), we manually define the ground-truth properties.
@@ IV-C1 Methodology
IV-C2 Results
Table II: Comparison between properties automatically extracted by our proposed approach and ground-truth properties that are manually crafted.
@@ IV-C2 Results
Pr9 Laser Tripwire “INTRUDER” is printed if an only if there is no light. Same as ground-truth.
@@
Pr14 Ultrasonic Thermin Volume level never gets too high or too low. Same as ground-truth.
Pr15 Ultrasonic Thermin Volume increases when user is getting closer to the sensor. Same as ground-truth.
Pr16 Ultrasonic Thermin Correctly not included. The buzzer has a correct range of volume. This is an irrelevant property that mostly checks the inner classes of the gpiozero library, not the application.
@@
Pr20 Presence Indicator The LEDs show the number of present people divided by 10. Same as ground-truth.
Pr21 Presence Indicator When the number of present people is above 10, the number “1” is shown. Same as ground-truth.
Pr22 Presence Indicator Correctly not included. When the PresneceIndicator object is closed, the LEDs are off. This is not a useful property, as it only tests very detailed implementation details.
Pr23 gpiozero InputDevice The pull_up parameter of the input device affects the pull state of the pin. Same as ground-truth.
Pr24 gpiozero InputDevice With a none pull_up, the active_state should be set. Same as ground-truth.
Pr25 gpiozero InputDevice When pull_up is set, the is_active parameter has a reverse effect between the pull_up value of the device and the activation of the pin. Same as ground-truth.
Pr26 gpiozero InputDevice Neglected. When the input device is closed, the pin is not in use anymore.
Table II shows the results of this experiment. In total, the table contains 26 properties. We split these properties into four groups: Group1 consists of 15 properties that are present among ground-truth and ChekProp extracted properties in the exact same form (Pr3, Pr5, Pr7, Pr9, Pr10, Pr11, Pr12, Pr14, Pr15, Pr17, Pr20, Pr21, Pr23, Pr24, and Pr25); Group2 consists of 3 properties that are present among ground-truth and ChekProp extracted properties in equivalent but slightly different forms (Pr1, Pr2, and Pr6); Group3 consists of 7 properties that are only among the ChekProp extracted properties (Pr4, Pr8, Pr13, Pr16, Pr19, Pr22, and Pr26); and Group4 consists of 1 property that is only among ground-truth properties (Pr18).
In Table II, we see that in three cases (Pr4, Pr8 and Pr26) ChekProp extracts a property that is relevant and useful, but neglected in manually crafted properties. For example, our approach extracts Pr4 for PCS which notes that neither heater or cooler should be active when the room temperature is between 21°C and 23°C. This indicates that not only can our automated approach replace manual property design work, but it can also even improve the manually crafted properties.
@@
As presented in Table II, four of the properties extracted by ChekProp (Pr13, Pr16, Pr19, and Pr22) are not useful. These properties either check a state that does not occur in real-wrold (Pr13) or validate highly detailed implementation nuances. This observation shows that a manual check on properties automatically extracted by ChekProp is needed to ensure that no useful property is considered for testing.
Finally, there is only one ground-truth property (Pr18) that does not correspond to any of the properties extracted by ChekProp. Pr18 is a property for the quick reaction game and indicates the order of changes in the light status, it should be first turned on at some point and then turned off at some point. With a careful manual analysis, we understand that the code we provide to the LLM for the quick reaction game lacks the documentation regarding this point. This suggests the importance natural language description as of one core components in ChekProp prompts (see Section III-A).
@@
IV-D RQ2: PBT Quality
IV-D1 Methodology
For evaluating the quality of PBTs generated by ChekProp, we examine the PBTs that test the 21 relevant properties according to our analysis in the RQ1 experiment (see Section IV-C). As explained in Section IV-A, we assess the applicability of our approach from two aspects: executability and effectiveness.
@@ IV-D1 Methodology
To assess the effectiveness of generated PBTs, we study if it checks the property over representatives of all or most partitions of the input space. As explained in Section II, one of the main components of a PBT is an input generator that produces various inputs from the input space. In this experiment, we determine to what extent the input generators of generated PBTs produce inputs from all partitions of the input space. The more partitions of input space considered by a PBT, the more effective the PBT is. We assess the effectiveness of PBTs generated for each program through a manual analysis and assign them to one of the three effective groups “HIGH”, “MED”, and “LOW”.
IV-D2 Results
Table III: The quality of PBTs generated by ChekProp for the 21 relevant propertie. A “HIGH”, “MED”, or “LOW” level in the “Executability” column indicates the PBT can be successfully executed with less than one minute, between one to three minutes, or more than three minutes of manual effort for modification, respectively. The “Effectiveness” column indicates the level of input space partitions covered by the PBT.
@@ IV-D2 Results
Pr20 Presence Indicator MED The used mocking method does not work for parameterized tests. MED
Pr21 Presence Indicator MED The used mocking method does not work for parameterized tests. MED
Pr23 gpiozero InputDevice MED The used mocking method does not work for parameterized tests. HIGH
Pr24 gpiozero InputDevice MED The used mocking method does not work for parameterized tests. HIGH
Pr25 gpiozero InputDevice LOW The used mocking method does not work for parameterized tests. The pin state is set with an incorrect use of the interface. HIGH
Pr26 gpiozero InputDevice MED The used mocking method does not work for parameterized tests. HIGH
Table III summarizes the result of our experiment on ChekProp applicability. The “Property_ID” and “Program” columns indicate the property and the program that the PBT is testing. Note that the ID of the property is taken from Table II, which lists the properties extracted by ChekProp. The “Executability” column shows the result of our assessment of the executability of generated PBTs in terms of their syntactical and semantical correctness. The fourth column presents the main executability challenge of generated PBTs that should be addressed manually. Finally, the last column presents the level of effectiveness of PBTs generated for each program.

### v2

@@ Abstract
Keywords:
Property-based Testing LLM4SE Cyber-Physical System Safety
1 Introduction
Cyber-physical systems (CPS) are integrated hardware-software systems where computation and physical processes are deeply intertwined. Ensuring safety [7] in these systems, in particular for the safety-critical ones is of high importance, as failures can have critical consequences. One of the key strategies in safety assurance is to capture the system’s properties describing what the system should and should not do under different conditions. Encoding the requirements as formal or semi-formal properties, can enable creating safety and security guardrails for system behavior. Formulating properties usually starts from the system requirements typically written in natural language. Therefore, large language models (LLMs) with their strong potential can be leveraged to extract properties from existing documentation and software code. These properties can be used to drive subsequent automated testing and verification activities, such as property-based testing [5, 3]. Property-based tests (PBT) are software tests that check a given property regarding the expected behavior holds for various input scenarios.
In this paper, we propose a novel automated and scalable approach for guardrailing CPS using PBT generated by LLMs. Our approach benefits from two major established facts in software engineering: 1) the cyber side of CPSs is essentially a software program amenable to existing automated program analysis tools; and 2) advanced LLMs are strong in analyzing programs and extracting their expected properties [20]. Based on these two observation, our proposed approach, called ChekProp, uses LLMs to generate property-based tests for CPSs before their deployment, i.e., at design time. These PBTs can also then be utilized after deployment, i.e., at run time, to detect unsafe behavior of the CPS. ChekProp uses the source code, documentation, and unit tests of the target CPS to extract properties regarding its expected behavior. ChekProp also generates PBTs that verify that the extracted properties hold for the CPS. We implement a prototype of ChekProp and make it publicly available to the community [4].
@@ 1 Introduction
2 Background on Property-based Testing
@@ 3.2 Initial PBT Generation
Note that, as explained in section 2, there is significant difference between unit tests and property based tests. The unit test only checks the behavior of the program for a specific input. For example, in the unit test (Section 3) of Figure 3, specific total_time, cylinder_interval, etc. are used. Also, in unit testing we usually check that the output is exactly what is expected according to an oracle [19]. In contrast, property based tests check that a more general condition is meet by the program behavior over a wide range of inputs.
The fourth and final part of the Figure 3 instructs the LLM to generate the desired PBTs. At the end of the initial PBT generation step, ChekProp obtains a set of initial PBTs. ChekProp runs these PBTs and collects their results using an analyzer unit. If a group of generated PBTs fails, ChekProp collects their failure message and enters its PBT improvement loop step as described in subsection 3.3.
@@
3.3 PBT Improvement Loop
@@ 3.5 Implementation
3.6 Property-based Monitoring
@@
More generally, as explained in section 2 and as seen in 2, the PBTs generated in the first phase consist of three components: an input generator, a test body, and a test assertion. In the property-based monitoring phase, the PBT input generator is no longer needed, as the inputs are generated by the cyber part (in the form of control commands) and the physical system (via sensor values). The test body is replaced by the monitor that sits between the controller and the physical system collecting relevant data. The properties derived at design time are transformed into guards that are checked at runtime. The monitor verifies that these guards hold in the current state of the system. Based on the situation, if a violation is detected, the monitor can intercept and block the command being sent to the physical system. In this way, the PBTs generated by ChekProp serve as runtime guardrails for CPSs.
4 Experiments
4.1 Research Questions
@@
4.2 Dataset
Table 1: The cyber-physical system Python programs considered in our dataset.
@@ 4.2 Dataset
P1 Temperature Control
System (TCS)
Model-based
Example [12]
@@
The TCS system controls the temperature of a room. The room has a window that blows cold or hot wind into the room. A sensor collects the current temperature and a HC unit is used to heat up or cool down the room environment. Finally, a controller collects the temperature data from the sensor and decides if the HC unit should be used.
P2 Pneumatic Control
System (PCS)
Model-based
Example [12]
@@
The PCS regulates the movement of components of a mechanical system that consist of a horizontal and a vertical cylinder, their corresponding controllers and sensors. This mechanical system is designed to pick an object from the ground and move it to a new place.
P3 Laser Tripwire Raspberry
Pi Projects [15]
@@
A prototype on a breadboard to detect whether a beam of light is hitting the light-dependent resistor (LDR).
P4 Line Following Robot Raspberry
Pi Projects [15]
@@
A robot with two motors and two sensors. The robot uses its right and left sensors to follow a line and decide if should go turn right, left, or go forward.
P5 Ultrasonic Theremin Raspberry
Pi Projects [15]
@@
A system that produces a sound with a volume level corresponding to the user distance from the device.
P6 Remote Buggy Raspberry
Pi Projects [15]
@@
A robot that moves in four directions and a user remotely controls it.
P7 Quick Reaction Game Raspberry
Pi Projects [15]
@@
A game in which a light turns on and off and two players compete to hit their button faster after the light is turned off.
P8 Presence Indicator Raspberry
Pi Projects [15]
@@
A monitor made of a list of LEDs that shows the number of people present at a place.
P9 gpiozero InputDevice gpiozero
Library Class[17]
@@
A core class in gpiozero library that handles connection to input devices on a Raspberry Pi board.
@@
As the current version of ChekProp supports PBT generation for Python programs, we curate a dataset of Python CPS programs for our experiments. This dataset consists of nine programs presented in Table 1. These programs are taken from three main sources as follows.
@@
First, we include the Python version of two CPSs that are widely studied in the model checking literature [12]: a temperature control system (TCS) and a pneumatic control system (PCS) (P1 and P2 in Table 1). TCS and PCS are presented with the IDs P1 and P2 in Table 1. Moradi et al. [12] use the model checking tool of Rebeca  [^1] Rebeca is an actor-based modeling language, in which actors are the units of concurrency and communicate via asynchronous message passing [16]., Afra [16], to detect potential attacks against these systems. For this, they define the correctness properties for each system and assess if Afra can find counterexamples for these properties on models augmented with malicious behavior. The manually defined properties in [12] set a ground-truth with which we can compare the properties automatically extracted by ChekProp.
The second set of programs in our dataset are the six programs (P3-P8 in Table 1) taken from open-source Raspberry Pi projects [15] that use the gpiozero library [17]. gpiozero is a Python library that real-world cyber-physical systems employ to connect to Raspberry Pi boards. We take the source of these programs from the official Raspberry Pi website [15] and manually add unit tests for them. These unit tests can be used in ChekProp prompts (see subsection 3.1). The six programs in our dataset that use gpiozero enable us to evaluate ChekProp applicability on projects that adopt the widely used Raspberry Pi boards.
@@ 69 If :data:‘True‘, the device uses a pull-up resistor to set the GPIO pin
The last program considered in our dataset (P9 in Table 1) is InputDevice, a core class from the gpiozero library. This class provides an interface for Python programs to interact with Raspberry Pi input devices, such as barometers, temperature sensors, etc. The InputDevice class is a complex class from gpiozero that connects to various components in this library. Testing this class requires a detailed understanding of the inner workings of the library and how it represents and handles the physical environment. For example, to instantiate an object of the InputDevice class, we have to pass the number of a pin to its constructor method. The type of this pin should be “input”, while some of the pins on a Raspberry Pi board are reserved only for “output”. A correct test should use a pin of the correct type to instantiate an InputDevice. Another example of such details in gpiozero is how pins are activated. To activate a pin on a Raspberry Pi board, its voltage should go high, which can happen by calling the pin.drive_high() method. Writing a test for a CPS program requires an accurate understanding of how interacting with the program, such as calling pin.drive_high(), impacts the physical status of the system, such as increasing the voltage on the Raspberry Pi board. To evaluate whether ChekProp generated PBTs capture such details about CPS programs, we include the InputDevice class in our dataset.
4 presents the InputDevice class. This class, similar to other gpiozero classes, has well-written documentation (lines #2-39). We consider this documentation as the natural language description in ChekProp prompts (see subsection 3.1). Moreover, gpiozero has extensive unit tests for its classes, which we use them to produce our initial prompt . This confirms that gpiozero classes have the essential components for applying ChekProp: the natural language description, the source code, and the unit test (see subsection 3.1).
@@
4.3 RQ1: Property Relevance
4.3.1 I. Methodology:
To assess the quality of extracted properties, we run our approach on the nine programs in our dataset and analyze the relevance of the extracted properties. In this experiment, we abstract away the implementation details of the generated PBTs; instead, we focus on how well the properties considered in the PBTs validate the logic of the program under test. For this purpose, we compare properties extracted by our approach with manually crafted properties that we consider as ground-truth. In particular, we evaluate whether the logic checked by ground-truth properties is also validated by ChekProp extracted properties and vice versa. As explained in subsection 4.2, for programs P1 and P2, the ground-truth properties are already stated by Moradi et al. [12]. For the remaining programs (P3-P9), we manually define the ground-truth properties.
@@ 4.3 RQ1: Property Relevance
4.3.2 II. Results:
LABEL:tab:rq1_res shows the results of this experiment. In total, the table contains 26 properties. We split these properties into four groups: Group1 consists of 15 properties that are present among ground-truth and ChekProp extracted properties in the exact same form ( Pr3, Pr5, Pr7, Pr9, Pr10, Pr11, Pr12, Pr14, Pr15, Pr17, Pr20, Pr21, Pr23, Pr24, and Pr25); Group2 consists of 3 properties that are present among ground-truth and ChekProp extracted properties in equivalent but slightly different forms (Pr1, Pr2, and Pr6); Group3 consists of 7 properties that are only among the ChekProp extracted properties (Pr4, Pr8, Pr13, Pr16, Pr19, Pr22, and Pr26); and Group4 consists of 1 property that is only among ground-truth properties (Pr18).
Table 2: Comparison between properties automatically extracted by our proposed approach and ground-truth properties that are manually crafted.
@@
Pr9 Laser Tripwire "INTRUDER" is printed if an only if there is no light. Same as ground-truth.
@@
Pr14 Ultrasonic
Thermin
Volume level never gets too high or too low. Same as ground-truth.
Pr15 Ultrasonic
Thermin
Volume increases when user is getting closer to the sensor. Same as ground-truth.
Pr16 Ultrasonic
Thermin
Correctly not included. The buzzer has a correct range of volume.This is an irrelevant property that mostly checks the inner classes of the gpiozero library, not the application.
@@
Pr20 Presence
Indicator
The LEDs show the number of present people divided by 10. Same as ground-truth.
Pr21 Presence
Indicator
When the number of present people is above 10, the number “1” is shown. Same as ground-truth.
Pr22 Presence
Indicator
Correctly not included. When the PresneceIndicator object is closed, the LEDs are off.This is not a useful property, as it only tests very detailed implementation details.
Pr23 gpiozero
InputDevice
The pull_up parameter of the input device affects the pull state of the pin. Same as ground-truth.
Pr24 gpiozero
InputDevice
With a none pull_up, the active_state should be set. Same as ground-truth.
Pr25 gpiozero
InputDevice
When pull_up is set, the is_active parameter has a reverse effect between the pull_up value of the device and the activation of the pin. Same as ground-truth.
Pr26 gpiozero
InputDevice
Neglected. When the input device is closed, the pin is not in use anymore.
In LABEL:tab:rq1_res, we see that in three cases (Pr4, Pr8 and Pr26) ChekProp extracts a property that is relevant and useful, but neglected in manually crafted properties. For example, our approach extracts Pr4 for PCS which notes that neither heater or cooler should be active when the room temperature is between 21°C and 23°C. This indicates that not only can our automated approach replace manual property design work, but it can also even improve the manually crafted properties.
@@
As presented in LABEL:tab:rq1_res, four of the properties extracted by ChekProp (Pr13, Pr16, Pr19, and Pr22) are not useful. These properties either check a state that does not occur in real-wrold (Pr13) or validate highly detailed implementation nuances. This observation shows that a manual check on properties automatically extracted by ChekProp is needed to ensure that no useful property is considered for testing.
Finally, there is only one ground-truth property (Pr18) that does not correspond to any of the properties extracted by ChekProp. Pr18 is a property for the quick reaction game and indicates the order of changes in the light status, it should be first turned on at some point and then turned off at some point. With a careful manual analysis, we understand that the code we provide to the LLM for the quick reaction game lacks the documentation regarding this point. This suggests the importance natural language description as of one core components in ChekProp prompts (see subsection 3.1).
@@
4.4 RQ2: PBT Quality
4.4.1 I. Methodology:
For evaluating the quality of PBTs generated by ChekProp, we examine the PBTs that test the 21 relevant properties according to our analysis in the RQ1 experiment (see subsection 4.3). As explained in subsection 4.1, we assess the applicability of our approach from two aspects: executability and effectiveness.
@@ 4.4 RQ2: PBT Quality
To assess the effectiveness of generated PBTs, we study if it checks the property over representatives of all or most partitions of the input space. As explained in section 2, one of the main components of a PBT is an input generator that produces various inputs from the input space. In this experiment, we determine to what extent the input generators of generated PBTs produce inputs from all partitions of the input space. The more partitions of input space considered by a PBT, the more effective the PBT is. We assess the effectiveness of PBTs generated for each program through a manual analysis and assign them to one of the three effective groups “HIGH”, “MED”, and “LOW”.
4.4.2 II. Results:
LABEL:tab:rq2_res summarizes the result of our experiment on ChekProp applicability. The “Property_ID” and “Program” columns indicate the property and the program that the PBT is testing. Note that the ID of the property is taken from LABEL:tab:rq1_res, which lists the properties extracted by ChekProp. The “Executability” column shows the result of our assessment of the executability of generated PBTs in terms of their syntactical and semantical correctness. The fourth column presents the main executability challenge of generated PBTs that should be addressed manually. Finally, the last column presents the level of effectiveness of PBTs generated for each program.
Table 3: The quality of PBTs generated by ChekProp for the 21 relevant propertie. A "HIGH", "MED", or "LOW" level in the "Executability" column indicates the PBT can be successfully executed with less than one minute, between one to three minutes, or more than three minutes of manual effort for modification, respectively. The "Effectiveness" column indicates the level of input space partitions covered by the PBT.
@@
Pr20 Presence
Indicator
MED The used mocking method does not work for parameterized tests. MED
Pr21 Presence
Indicator
MED The used mocking method does not work for parameterized tests. MED
Pr23 gpiozero
InputDevice
MED The used mocking method does not work for parameterized tests. HIGH
Pr24 gpiozero
InputDevice
MED The used mocking method does not work for parameterized tests. HIGH
Pr25 gpiozero
InputDevice
LOW The used mocking method does not work for parameterized tests.
The pin state is set with an incorrect use of the interface.
HIGH
Pr26 gpiozero
InputDevice
MED The used mocking method does not work for parameterized tests. HIGH